OpenStudy (anonymous):

Help D: 4x - 2x - 2 + -42 3n + 1 = 16 0.25(7 + 3m) = -0.125m SHOW WORK ;o

OpenStudy (anonymous):

What are you trying to solve?

OpenStudy (anonymous):

x , n and m

OpenStudy (anonymous):

ohh okay, give me a sec

OpenStudy (anonymous):

for the first problem, make sure you copied it correctly

OpenStudy (anonymous):

do u mean 4x - 2x - 2 = -42?

OpenStudy (anonymous):

haa lol i meant =

OpenStudy (anonymous):

kk wat did u get as ur answer for the first one?

OpenStudy (anonymous):

for the second one N=5

OpenStudy (anonymous):

subtract 1 from both sides it gives you 3n=15 then divide by 3 and you get n=5

OpenStudy (anonymous):

i cant get the first one. - hawk.

OpenStudy (anonymous):

4x - 2x - 2 = -42 2x - 2 = -42 2x = -40 x = -20

OpenStudy (anonymous):

kevin let her try the last one for herself

OpenStudy (anonymous):

we'll correct it for u though :)

OpenStudy (anonymous):

alright haha

OpenStudy (anonymous):

so chubby bunny, do u understand how we both solved for x and n?

OpenStudy (anonymous):

haha yeah. and the last one i got -40.5 ;o

OpenStudy (anonymous):

you guys don mind that i have one more ?

OpenStudy (anonymous):

ok i got something different

OpenStudy (anonymous):

i dont :)

OpenStudy (anonymous):

0.25(7 + 3m) = -0.125m 7 + 3m = -0.5m U got that?

OpenStudy (anonymous):

i got it now :)

OpenStudy (anonymous):

so wats ur answer now?

OpenStudy (anonymous):

2 ? /.\

OpenStudy (anonymous):

i got m = -2.5

OpenStudy (anonymous):

Yes, its -2.5

OpenStudy (anonymous):

oh crap lol i mad ea mis take xD

OpenStudy (anonymous):

Haha, i was about to tell you that hawk

OpenStudy (anonymous):

@ChubbyBunny~ use a calculator for the decimal ones

OpenStudy (anonymous):

ok ill show full working now

OpenStudy (anonymous):

i was going to , but i don have one ;/

OpenStudy (anonymous):

7 + 3m = -0.5m 3.5 m = -7 m = -2

OpenStudy (anonymous):

use the google one, just google calculator it will show up on your screen

OpenStudy (anonymous):

so u were close C.Bunny :)

OpenStudy (anonymous):

it is -2 not 2

OpenStudy (anonymous):

google has everything ... - kevin

OpenStudy (anonymous):

haa , then good :) - hawk

OpenStudy (anonymous):

@HawkCrimson dont you multiply 0.25 by 7 and 3?

OpenStudy (anonymous):

back, wat r ur other questions :)

OpenStudy (anonymous):

or u dont need any more help?

OpenStudy (anonymous):

did u understand how i solved for m?

OpenStudy (anonymous):

i still need hep ; _ ; yuss , i understood ~

OpenStudy (anonymous):

ok wat r the other questions? :) or do u want me to give u some?

OpenStudy (anonymous):

a/4 - 9 = 5 , lol what do you mean ?

OpenStudy (anonymous):

i can make up some for u to practice :) but if u have other questions then that will be better

OpenStudy (anonymous):

were u able to get an answer for a?

OpenStudy (anonymous):

oh lol practice would be good ;D i suck at math ... ; - ;

OpenStudy (anonymous):

not if i can help it ;)

OpenStudy (anonymous):

nu uh .. i dont know what a is D:

OpenStudy (anonymous):

haa , right ? c;

OpenStudy (anonymous):

Ill start u off :) a/4 - 9 = 5 a/4 = 14

OpenStudy (anonymous):

can u solve for a rom there?

OpenStudy (anonymous):

@HawkCrimson hawk do you know geometry?

OpenStudy (anonymous):

you multiply 4 and 14 , right ?

OpenStudy (anonymous):

yesh! :D

OpenStudy (anonymous):

its 56 ?? (:

OpenStudy (anonymous):

yay! congrates ;D

OpenStudy (anonymous):

ill be here whole day so u can just ask as much questions as u would like and ill help :)

OpenStudy (anonymous):

if i leave then the site will notify me and ill return ;D

OpenStudy (anonymous):

okay ^ - ^ message me ?

OpenStudy (anonymous):

sure :)